OpenStudy (anonymous):

let lim of f(x)=100 find lim f sq rt. f(x) lim as x approaches -9 a)10 b)100 c)3.1623 d)-9

hartnn (hartnn):

is it \(\large f\sqrt {f(x)}\) ?? or just sqrt f(x) ?

OpenStudy (anonymous):

bottom is correct

hartnn (hartnn):

so just \(\sqrt {f(x)}\) since f(x) = 100 can you find \(\sqrt {f(x)}\) ??

OpenStudy (anonymous):

10

OpenStudy (anonymous):

thanks

hartnn (hartnn):

yes, limit of constant is that constant itself, wherever the limit tends to :)